OpenStudy (anonymous):

(x + 20) + (2x + 10) = 90 57 40 30

OpenStudy (anonymous):

x=20 I think your answers are wrong.

jhonyy9 (jhonyy9):

3x=90-30 3x=60 x=60/3 x=20